HIGH-RESOLUTION LASER-INDUCED FLUORESCENCE SPECTROSCOPY OF CYCLOHEXOXY: RO-TATIONAL AND FINE STRUCTURE OF MOLECULES IN NEARLY DEGENERATE ELECTRONIC STATES

\underline{JINJUN LIU}, Department of Chemistry, University of Louisville, 2320 S. Brook St., Louisville, Kentucky 40292.; DMITRY G. MELNIK, and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Ave., Columbus, Ohio 43210.

The previously obtained B̃^{2}A′ −X̃^{2}A′′ and B̃^{2}A′ −Ã^{2}A′ laser-induced fluorescence (LIF) spectra of jet-cooled cyclohexoxy radical (c-C_{6}H_{11}O)^{a} have been analyzed and simulated using the coupled-two-state model presented in the preceding talk. The rotational and fine structure of the nearly degenerate X̃^{2}A′′ and Ã^{2}A′ states is reproduced using one set of molecular constants including rotational constants, spin-rotation constants, effective spin-orbit constants (aζ_{e}d) and the vibronic energy separation between the two states (ΔE). While the energy level structure could be reproduced by only effective spin-rotation constants (without the spin-orbit constant), the spin-orbit interaction introduces transitions that have no intensity using the separate-states asymmetric rotor model. Rotational and fine-structure analysis using the two-state model has proven to be an effective method to separate the first order electron-spin-molecular-rotation constants from the effective spin-rotation constants, and to decouple the spin-orbit splitting (aζ_{e}d) and the vibronic energy separation (ΔE), both of which contribute to the experimentally observed energy separation between the two coupled states (ΔE^{Ã−X̃}). Isopropoxy (discussed in the preceding talk), cyclohexoxy, and other molecules in nearly degenerate electronic states provide unique cases bridging the gap from symmetrically degenerate states, e.g., ground X̃^{2}E state of methoxy, and the Born-Oppenheimer limit of unperturbed electronic states.

a“Jet-cooled laser spectroscopy of the cyclohexoxy radical”, L. Zu, J. Liu, G. Tarczay, P. Dupré, and T. A. Miller, J. Chem.

Phys. 120, 10579 (2004).